How many Suns would it take, laid side by side, to reach a star 5.0 light-years away?
Answer #1

Diameter of sun = 1.392 x 106 Km

= 1.392 x 109 m

One light year is distance equivalent to the distance that light travels in one year.

One light year = 9.4607 × 1015 m

5 light year = 47.3 × 1015 m

numbers of sun = 5 light year distance / diameter of one sun

= 47.3 × 1015 / 1.392 x 109

numbers of sun = 33.98 x 106